# Octopus: Experiences with a Hybrid Event-Driven Architecture for Distributed Scientific Computing

### v1

@@ Abstract
Scientific research increasingly relies on distributed computational resources, storage systems, networks, and instruments, ranging from HPC and cloud systems to edge devices. Event-driven architecture (EDA) benefits applications targeting distributed research infrastructures by enabling the organization, communication, processing, reliability, and security of events generated from many sources. To support the development of scientific EDA, we introduce Octopus, a hybrid, cloud-to-edge event fabric designed to link many local event producers and consumers with cloud-hosted brokers. Octopus can be scaled to meet demand, permits the deployment of highly available Triggers for automatic event processing, and enforces fine-grained access control. We identify requirements in self-driving laboratories, scientific data automation, online task scheduling, epidemic modeling, and dynamic workflow management use cases, and present results demonstrating Octopus’ ability to meet those requirements. Octopus supports producing and consuming events at a rate of over 4.2 M and 9.6 M events per second, respectively, from distributed clients.
@@ I Introduction
The development of such applications is complicated by the need simultaneously to address familiar distributed computing concerns such as failure while also meeting challenging performance requirements and dealing with the unique properties of scientific infrastructure. Methods used historically to develop scientific applications within single machines or institutions, like message passing and workflow tools, fare badly on dynamic, heterogeneous DRI, in which anomalies are routine rather than unexpected. At the same time, the event-driven architectures (EDA) used to construct reliable and scalable distributed applications [11, 12, 13] struggle with the often large data volumes of scientific applications, and with specialized properties of DRI, from unique event sources and event handling actions to large number of authorization domains and the overhead to provide robust and scalable multi-user capabilities on which scientists can depend.
@@
As part of our investigation of these questions, we have developed Octopus, [^1] Name anonymized for double-blind review. a prototype global event fabric and software ecosystem that enables scientific users and applications to communicate and process events within DRI. We design Octopus around a globally-accessible, cloud-hosted, multi-user service with fine-grained access control, enabling the secure publication and consumption of events from any internet-accessible location: see Figure 1.
To address application requirements, we implement functionality in Octopus to process events and to trigger arbitrary actions (e.g., via web services) across the distributed ecosystem. Octopus users can define rules that specify the actions to take when particular events are received. For example, a first rule might state that data acquisition at an instrument should trigger a workflow to transfer the data to an HPC system; a second that completion of the transfer should trigger analysis on the HPC; and a third that conclusion of the analysis should trigger an email to a researcher with results.
We are concerned with both the performance and practical usability of Octopus. To address the first question, we present microbenchmarks that quantify its scalability and performance. To investigate the second, we examine its use in five applications; a self-driving laboratory, distributed data monitoring and automation, FaaS scheduling, epidemic modeling, and workflow monitoring. In each case, we show how Octopus, and more generally, EDA, can enable new functionality not previously possible.
@@
    The proposal of scientific EDA to design applications and a survey of use cases to determine requirements.
@@
    The design and implementation of a hybrid, cloud-hosted event fabric, Octopus, and its software ecosystem.
@@
    The evaluation of Octopus’ performance, scalability, and suitability for scientific use cases.
  4. 4.
    The implementation of triggers to filter and act on events for real-time, reliable, and automatic responsiveness.
  5. 5.
    Discussion of the benefits and limitations of applying EDA and Octopus to five science applications.
@@ III Use Cases & Requirements
SDL $\sim$100×$\times R$ 0.5 KB 1 $R$ 1
Data Auto. $\sim$1,000×$\times R$ 4 KB 1 $R$ Trigger
Scheduling $\sim$10,000×$\times R$ 1 KB $R$ $R$ 1
Epidemic $\sim$10101010×$\times R$ 1 KB $R$ $R$ Trigger
Workflow $\sim$5,000×$\times R$ 1 KB $R$ $R$ $R$
@@ IV Design and Implementation
Figure 2: Octopus architecture. Users interact with the web service (green arrows) to acquire credentials, where their identity is mapped to an IAM entity and shared with the MSK ZooKeeper. Producers (left) and consumers (right) communicate with the event fabric to publish and receive events, respectively (blue arrows). A trigger can be configured to act on events and invoke remote actions (yellow arrows). Events can also be persisted to reliable cloud storage when enabled (red arrows).
@@ IV-A Event Fabric
We build the Octopus event fabric on Apache Kafka, an event streaming platform with high throughput, reliability, and scalability [25]. A Kafka deployment runs on a cluster of one or more servers called brokers, which expose a publish/consume interface and handle the receipt of messages from producers and their dispatch to consumers. Messages are grouped first into topics and then partitions, with messages within a partition strictly ordered. Large Kafka clusters, with more than $4000$ brokers, have been used to process over 7 trillion messages daily [26].
Rather than host our own Kafka cluster, we use AWS Managed Streaming for Kafka (MSK)—a highly available, scalable, and robust Kafka deployment that benefits from integration with other AWS services, such as IAM identity management, S3 for persistent storage, and Amazon CloudWatch for monitoring. MSK allows the Octopus event fabric to be scaled dynamically and geo-replicated for reliability across AWS service regions, while enforcing fine-grained access control that can restrict access on a per-topic level for authenticated users. One potential downside, discussed below, are the costs associated with this use of AWS services.
@@ IV-B Octopus Web Service
The Octopus Web Service (OWS) allows users to provision, configure, and manage topics; acquire and manage credentials to access MSK topics; and deploy and manage triggers. OWS is implemented as a RESTful service deployed on AWS EC2 instances. OWS implements a comprehensive security model, using OAuth 2.0, via which users must first authenticate via Globus Auth [27], and include access tokens in subsequent API requests. OWS exposes REST routes to provision, configure, share, or release topics:
@@
  • POST <service>/topic/<topic>/partitions: Set the number of partitions for topic.
@@ IV-C Authentication and Authorization
Authentication and authorization of event producers is important for provenance and accountability, and of event consumers to protect sensitive information. We build our authentication and authorization model on Globus Auth, as it: 1) is a standards-compliant OAuth 2.0 implementation; 2) supports a wide range of identity providers, including many commonly used in research; 3) implements a unique delegation model via which users can delegate permission for services to call other services (e.g., for an external service to register an event in Octopus, or for Octopus triggers to call external actions, on behalf of a user); and 4) is used by many science services, simplifying their integration as actions in Octopus.
@@ IV-D Octopus Triggers
We employ the function-as-a-service (FaaS) paradigm to implement a managed trigger model with these capabilities. Specifically, we deploy, for each trigger, an AWS Lambda function to process events from the associated topic and invoke specified actions. Users can specify the Lambda function and execution environment for that function. We use AWS EventBridge event patterns [30] to perform optional event filtering to trigger the Lambda function only on events of interest. EventBridge patterns use a simple JSON-based language for specifying rules (e.g., see 1). When a user registers a trigger, OWS configures the optional pattern and also creates the appropriate IAM policy, IAM role, and CloudWatch log group to manage and monitor the Lambda function.
@@
Each Lambda function is given its own MSK consumer group, meaning that many instances of the Lambda function can retrieve events without affecting other consumers of the topic. Lambda functions can scale automatically by evaluating processing pressure (the number of pending events in a topic). Lambda evaluates the processing pressure at 1 min intervals, and scales concurrent invocations of the function dynamically when warranted. The user can configure (via Octopus) the function to process batches of up to $10000$ events (or a total of 6 MB) per invocation. Triggers can be managed by:
@@ IV-E Octopus SDK
We provide a Python SDK to interact with the Octopus Web Service and simplify integration with existing applications and services. The SDK is available on PyPI and includes a Globus Auth login manager to perform an authentication flow and cache tokens on the user’s behalf. Tokens and MSK secrets are stored in a local SQLite database and automatically refreshed as needed. Optionally, the SDK can install kafka-python and aws-msk-iam-sasl-signer-python packages such that clients can use IAM authentication to interact with the event fabric and produce/consume events. Additionally, other Kafka producer/consumer toolkits, such as the AWS-provided Java toolkit, that support IAM authentication can also use these stored tokens to interact with the Octopus cluster.
@@ IV-F Scalability and Fault-tolerance
Each Octopus component is designed to be independently and dynamically scalable, adapting to real-time demand and scheduled scaling activities. The MSK cluster can be scaled up by resizing cloud instances and by adding additional brokers for horizontal scale up, and fault tolerance can be improved by replicating the cluster across regions. Topics may be replicated and synchronized by using the Kafka MirrorMaker tool. Managed consumer functions dynamically scale according to the processing pressure, which is reevaluated each minute to apply the scaling decisions. Octopus Web Service, deployed on EC2 instances, can be horizontally scaled behind an elastic load balancer. API operations on the OWS side are programmed to be idempotent such that the automatic retry of the function would not cause the system to be in inconsistent states. The source of truth of about which topics are owned by which identities are stored in ZooKeeper and replicated to IAM, meaning that scaling OWS does not affect data consistently.
@@ V-A Experimental Setup
TABLE III: Baseline performance and scalability benchmarking results. Producer and consumer throughputs are events/sec; latency is milliseconds (ms). A ‘-’ signifies no change from the above result.
@@ V-B Data Collection Process
Figure 3: Median (top) and 99 percentile (bottom) latencies vs. throughput for configurations 1–6 on baseline cluster with remote producers.
@@ V-C Baseline Performance
Acknowledgments: Experiments #3 and 4 change the acknowledgment requirements from 0 to 1 (more precisely, the leading one), and finally, all brokers must acknowledge, while keeping the event size constant at 1 KB. The peak throughput drops from 195 K (acks=0) to 161 K (acks=1) down to 65 K (acks=all) event/s for local producers, and from 174 K to 143 K down to 65 K event/s for remote producers. The median latency increases by 9 ms and 101 ms as well as 16 ms and 62 ms for the local and remote producers.
@@ V-D Octopus Triggers
Figure 5: Producer and consumer throughput vs. number of topics. For producer, increases to 273 K $\times$ 1 KB event/s at 4 topics and then remains flat. For consumer, increases to 846 K $\times$ 1 KB event/s at 16 topics.
@@ VI-A Self-driving Laboratories
Here we tackle the challenge of monitoring the many different events generated by an SDL. [^2] Anonymized for double-blind review. The SDL uses Octopus to create a global log of distributed actions spanning robotic devices, HPC resources, and data resources. Events include the initiation of experiments, transitions between stages of the workflow, and the collection of results, across various instruments and computational resources. This log facilitates transparent and real-time insights into ongoing experiment workflows as well as enables researchers to trace back through the decision-making and experiment processes for review or optimization. Furthermore, this detailed provenance is invaluable for ensuring reproducibility by recording the lineage of data and will eventually serve as the basis for event-driven automation. The event log is consumed to provide graphical representations of the experiment, allowing administrators to quickly evaluate throughput, understand the stage of various workflows, and inspect event-level information.
Figure 6: Three scientific EDAs implemented for our use cases: file system synchronization (left), with file creation events from one FS triggering data replication requests (see Section VI-B); epidemic modeling (middle), with events from data sources triggering model inference on new data (see Section VI-D); and online task scheduling (right), with events carrying energy information to guide scheduling decisions (see Section VI-E).
@@ VI-D Epidemic Modeling and Response
Researchers have adopted an EDA for a real-time epidemic monitoring and response platform [36]. This system monitors various web-based data sources (e.g., public health data), and when data are updated, it ingests, cleans, and validates the data. Prediction models are regularly retrained and run, and data and model results are published for decision makers. The platform uses Octopus to manage the various events that trigger modeling. Octopus triggers are utilized to execute cloud-hosted workflows for data ingestion and model execution. Events include timer-based events to retrieve updates periodically from the various data sources; detection of anomalies; data updates; and new model results.
@@ VII-A Scientific EDA Experiences
How can we build a multi-tenant EDA that serves the needs and meets the scale of broad science use cases? We have evaluated and used Octopus to implement several representative EDAs. Our approach, building on scalable cloud platforms and research services, provides confidence that Octopus can scale to meet the needs of diverse scientific applications with varying data, throughput, and latency requirements. The adoption of well-defined APIs ensures that Octopus can be extended to address future requirements.
@@ VIII Related work
The high data rates and inherent distribution of science necessitate event filtering and aggregation at the edge [49]. Hierarchical message passing has also proven useful in other contexts, such as federated learning communication [50] and provides an area of future work to determine how such practices should be applied in HPC settings.
Event-driven science: Researchers at the Large Hadron Collider (LHC) use an EDA to process over one million events per second [49], with the first filtering layer performing a 1/500 data reduction. The remaining events are then passed to other services for analysis, monitoring, and storage. These rates demonstrate the need for hierarchical filtering.
The Line Information System Architecture [51], LISA, is an EDA for industry 4.0 factory design and construction. LISA aims to increase flexibility and simplify the integration of industrial devices and services to establish control systems. Similarly, EUPHORIA [52] is designed to orchestrate communication between devices in a smart environment. This work reflects our findings that processing time is dependent on event size. Octopus differs from these applications by providing the infrastructure to implement arbitrary scientific EDAs with diverse requirements and focuses on establishing a global event management model using a hybrid cloud-edge architecture.

### v2

@@ Abstract
Scientific research increasingly relies on distributed computational resources, storage systems, networks, and instruments, ranging from HPC and cloud systems to edge devices. Event-driven architecture (EDA) benefits applications targeting distributed research infrastructures by enabling the organization, communication, processing, reliability, and security of events generated from many sources. To support the development of scientific EDA, we introduce Octopus, a hybrid, cloud-to-edge event fabric designed to link many local event producers and consumers with cloud-hosted brokers, and to provide a fabric for developing resilient applications. Octopus can be scaled to meet demand, permits the deployment of highly available Triggers for automatic event processing, and enforces fine-grained access control. We identify requirements in self-driving laboratories, scientific data automation, online task scheduling, epidemic modeling, and dynamic workflow management use cases, and present results demonstrating Octopus’ ability to meet those requirements. Octopus supports producing and consuming events at a rate of over 4.2 M and 9.6 M events per second, respectively, from distributed clients.
@@ I Introduction
The development of such applications is complicated by the need to address traditional distributed computing challenges, such as handling failures, while also meeting stringent performance requirements and accommodating the unique properties of scientific infrastructure. Methods used historically to develop scientific applications within single machines or institutions, like message passing and workflow tools, fare badly on dynamic, heterogeneous DRI, in which anomalies are routine rather than unexpected. At the same time, the event-driven architectures (EDA) used to construct reliable and scalable distributed applications [11, 12, 13] struggle with the often large data volumes of scientific applications, and with specialized properties of DRI, from unique event sources and event handling actions to large number of authorization domains and the overhead to provide robust and scalable multi-user capabilities on which scientists can depend.
@@
As part of our investigation of these questions, we have developed Octopus, a prototype global event fabric and software ecosystem that enables scientific users and applications to communicate and process events within DRI. We design Octopus around a globally-accessible, cloud-hosted, multi-user service with fine-grained access control, enabling the secure publication and consumption of events from any internet-accessible location: see Figure 1. To address application requirements, we implement functionality in Octopus to process events and to trigger arbitrary actions (e.g., via web services) across the distributed ecosystem. Octopus users can define rules that specify the actions to take when particular events are received. For example, a first rule might state that data acquisition at an instrument should trigger a workflow to transfer the data to an HPC system; a second that completion of the transfer should trigger analysis on the HPC; and a third that conclusion of the analysis should trigger an email to a researcher with results.
We are concerned with both the performance and practical usability of Octopus. To address the first question, we present microbenchmarks that quantify its scalability and performance. To investigate the second, we examine its use in five applications: a self-driving laboratory, distributed data monitoring and automation, FaaS scheduling, epidemic modeling, and workflow monitoring. In each case, we show how Octopus, and more generally, EDA, can enable new functionality not previously possible.
@@
    Investigation of using scientific EDA to design robust distributed applications, a survey of use cases to determine requirements, and discussion of the benefits and limitations of applying EDA to five science applications.
@@
    Design and implementation of a hybrid, cloud-hosted, multi-user event fabric, Octopus, along with its open-sourced software ecosystem. The Octopus SDK and a walkthrough notebook are available at github.com/globus-labs/diaspora-event-sdk.
@@
    Evaluation of Octopus’ performance, scalability, and suitability for scientific use cases. The evaluation methodology and results are detailed at doi.org/10.5281/zenodo.10975534
@@ III Use Cases & Requirements
SDL $O(10^{2})\times R$ 0.5 KB 1 $R$ 1
Data Auto. $O(10^{3})\times R$ 4 KB 1 $R$ Trigger
Scheduling $O(10^{4})\times R$ 1 KB $R$ $R$ 1
Epidemic $O(10)\times R$ 1 KB $R$ $R$ Trigger
Workflow $O(10^{3})\times R$ 1 KB $R$ $R$ $R$
@@ IV Design and Implementation
Figure 2: Octopus architecture. Users interact with the web service (green arrows) to acquire credentials, where their identity is mapped to an IAM entity and shared with the MSK ZooKeeper. Producers (left) and consumers (right) communicate with the event fabric to publish and receive events, respectively (blue arrows). A trigger can be configured to act on events and invoke remote actions (yellow arrows). Events can also be persisted to reliable cloud storage when enabled (red arrows). Monitoring consoles are available for admins to monitor the system’s live status and historical statistics.
@@ IV-A Event Fabric
We build the Octopus event fabric on Apache Kafka, an event streaming platform with high throughput, reliability, and scalability [25]. A Kafka deployment runs on a cluster of one or more servers called brokers, which expose a publish/consume interface and handle the receipt of messages from producers and their dispatch to consumers. Messages are grouped first into topics and then partitions, with messages within a partition strictly ordered. Large Kafka clusters, with more than $4000$ brokers, have been used to process over 7 trillion messages daily [26]. Empowered by Kafka, Octopus provides a substrate for both simple and complex workflows through general eventing interfaces, making it ideally suited for complex, real-time, asynchronous workflows.
Rather than host our own Kafka cluster, we use AWS Managed Streaming for Kafka (MSK)—a highly available, scalable, and robust Kafka deployment that benefits from integration with other AWS services, such as IAM identity management, S3 for persistent storage, and Amazon CloudWatch for monitoring. MSK allows the Octopus event fabric to be scaled dynamically and geo-replicated for reliability across AWS service regions, while enforcing fine-grained access control that can restrict access on a per-topic level for authenticated users. One potential downside, discussed below, is the costs associated with this use of AWS services.
@@ IV-B Octopus Web Service
The Octopus Web Service (OWS) allows users to provision, configure, and manage topics; acquire and manage credentials to access MSK topics; and deploy and manage triggers. OWS is implemented as a RESTful service deployed on AWS Lightsail instances. OWS implements a comprehensive security model, using OAuth 2.0, via which users must first authenticate via Globus Auth [27], and include access tokens in subsequent API requests. OWS exposes REST routes to provision, configure, share, or release topics:
@@
  • POST <service>/topic/<topic>/partitions: Set the number of partitions for the topic.
@@ IV-C Authentication and Authorization
Authentication and authorization of event producers are important for provenance and accountability and for event consumers to protect sensitive information. We build our authentication and authorization model on Globus Auth, as it: 1) is a standards-compliant OAuth 2.0 implementation; 2) supports a wide range of identity providers, including many commonly used in research; 3) implements a unique delegation model via which users can delegate permission for services to call other services (e.g., for an external service to register an event in Octopus, or for Octopus triggers to call external actions, on behalf of a user); and 4) is used by many science services, simplifying their integration as actions in Octopus.
@@ IV-D Octopus Triggers
We employ the function-as-a-service (FaaS) paradigm to implement a managed trigger model with these capabilities. This approach eliminates the need for users to define, deploy, or manage their own software for responding to events. Specifically, we deploy, for each trigger, an AWS Lambda function to process events from the associated topic and invoke specified actions. Users can specify the Lambda function and execution environment for that function. We use AWS EventBridge event patterns [30] to perform optional event filtering to trigger the Lambda function only on events of interest. EventBridge patterns use a simple JSON-based language for specifying rules (e.g., see 1) . When a user registers a trigger, OWS configures the optional pattern and also creates the appropriate IAM policy, IAM role, and CloudWatch log group to manage and monitor the Lambda function.
@@
Each Lambda function is given its own MSK consumer group, meaning that many instances of the Lambda function can retrieve events without affecting other consumers of the topic. Lambda functions can scale automatically by evaluating processing pressure (the number of pending events in a topic). Lambda evaluates the processing pressure at 1 min intervals, and scales concurrent invocations of the function dynamically when warranted. The user can configure (via OWS) the function to process batches of up to $10000$ events (or a total of 6 MB) per invocation. Octopus Triggers are versatile, secure, robust, and automated, providing a new way for developers to build scientific applications. Triggers can be managed by:
@@ IV-E Octopus SDK
We provide a Python SDK to interact with the Octopus Web Service and simplify integration with existing applications and services. The SDK is available on PyPI and includes a Globus Auth login manager to perform an authentication flow and cache tokens on the user’s behalf. Tokens and MSK secrets are stored in a local SQLite database and automatically refreshed as needed. Optionally, the SDK can install kafka-python to interact with the event fabric and produce/consume events. Additionally, other Kafka producer/consumer libraries that support IAM authentication can also use these stored tokens to interact with the Octopus cluster.
@@ IV-F Scalability and Fault-tolerance
Each Octopus component is designed to be independently and dynamically scalable, adapting to real-time demand and scheduled scaling activities. The MSK cluster can be scaled up by resizing cloud instances and by adding additional brokers for horizontal scale up, and fault tolerance can be improved by replicating the cluster across regions. Topics may be replicated and synchronized by using the Kafka MirrorMaker tool. Managed consumer functions dynamically scale according to the processing pressure, which is reevaluated each minute to apply the scaling decisions. Octopus Web Service, deployed on Lightsail instances, can be horizontally scaled behind an elastic load balancer. API operations on the OWS side are programmed to be idempotent such that the automatic retry of the function would not cause the system to be in inconsistent states. The source of truth about which topics are owned by which identities are stored in ZooKeeper and replicated to IAM, meaning that scaling OWS does not affect data consistently. Although ZooKeeper provides strong consistency for topic ownership, ownership updates are infrequent, so there is no bottleneck, and only authenticated users trigger updates through OWS.
@@ V-A Experimental Setup
TABLE III: Baseline performance and scalability benchmarking results. Producer and consumer throughputs are events/sec; latency is in milliseconds (ms). A ‘-’ signifies no change from the above result.
@@ V-B Data Collection Process
Figure 3: Median (top) and 99th percentile (bottom) latencies vs. throughput for configurations 1–6 on baseline cluster with remote producers.
@@ V-C Baseline Performance
Acknowledgments: Experiments #3 and #4 change the acknowledgment requirements from 0 to 1 (more precisely, the leading one), and finally, all brokers must acknowledge, while keeping the event size constant at 1 KB. The peak throughput drops from 195 K (acks=0) to 161 K (acks=1) down to 65 K (acks=all) event/s for local producers, and from 174 K to 143 K down to 65 K event/s for remote producers. The median latency increases by 9 ms and 101 ms as well as 16 ms and 62 ms for the local and remote producers.
@@ V-D Octopus Triggers
Figure 5: Producer and consumer throughput vs. number of topics. For producer, increases to 273 K $\times$ 1 KB event/s at four topics and then remains flat. For consumer, increases to 846 K $\times$ 1 KB event/s at 16 topics.
@@ VI-A Self-driving Laboratories
Here we tackle the challenge of monitoring the many different events generated by an SDL at Argonne National Laboratory. The SDL uses Octopus to create a global log of distributed actions spanning robotic devices, HPC resources, and data resources. Events include the initiation of experiments, transitions between stages of the workflow, and the collection of results, across various instruments and computational resources. This log facilitates transparent and real-time insights into ongoing experiment workflows as well as enables researchers to trace back through the decision-making and experiment processes for review or optimization. Furthermore, this detailed provenance is invaluable for ensuring reproducibility by recording the lineage of data and will eventually serve as the basis for event-driven automation. The event log is consumed to provide graphical representations of the experiment, allowing administrators to quickly evaluate throughput, understand the stage of various workflows, and inspect event-level information.
Figure 6: Three scientific EDAs implemented for our use cases: file system synchronization (left), with file creation events from one FS triggering data replication requests (see Section VI-B); online task scheduling (middle), with events carrying energy information to guide scheduling decisions (see Section VI-C); and epidemic modeling (right), with events from data sources triggering model inference on new data (see Section VI-D).
@@ VI-D Epidemic Modeling and Response
Researchers have adopted an EDA for a real-time epidemic monitoring and response platform [36, 37]. This system monitors various web-based data sources (e.g., public health data), and when data are updated, it ingests, cleans, and validates the data. Prediction models are regularly retrained and run, and data and model results are published for decision makers. The platform uses Octopus to manage the various events that trigger modeling. Octopus triggers are utilized to execute cloud-hosted workflows for data ingestion and model execution. Events include timer-based events to retrieve updates periodically from the various data sources; detection of anomalies; data updates; and new model results.
@@ VII-A Scientific EDA Experiences
How can we build a multi-tenant EDA that serves the needs and meets the scale of broad science use cases? We have evaluated and used Octopus to implement several representative EDAs in many environments, including Cloud (Scheduling), Edge (Epidemic), and HPC (Data Automation and Workflow). Our studies support our contributions, and our evaluations include configurations that represent large deployments of scientific applications. For example, we test with rates exceeding 10,000 events/second and 100 event consumers running at full capacity (Table III). These results are relevant for HPC clusters with thousands of computational nodes because Octopus is intended to be used through hierarchical aggregation (see Data Automation application), producing batched events sent to the cloud.
Our approach, building on scalable cloud platforms and research services, provides confidence that Octopus can scale to meet the needs of diverse scientific applications with varying data, throughput, and latency requirements. The adoption of well-defined APIs ensures that Octopus can be extended to address future requirements.
@@ VIII Related work
Event-driven science: The high data rates and inherent distribution of science necessitate event filtering and aggregation at the edge [50]. Researchers at the Large Hadron Collider (LHC) use an EDA to process over one million events per second [50], with the first filtering layer performing a 1/500 data reduction. The remaining events are then passed to other services for analysis, monitoring, and storage. These rates demonstrate the need for hierarchical filtering. Hierarchical filtering has also proven useful in other contexts, such as federated learning communication [51] and provides an area of future work to determine how such practices should be applied in HPC settings.